OpenStudy (anonymous):

Please help!!?? :) could some one go over this with me step by step so I can understand a little better. I already know the factors are 1,2,3,4,6,8,12,16,24,32,ext and that the final answer is 4sqrt6. Question: Simplify sqrt96

OpenStudy (raden):

use the property of sqrt(a * b) = sqrt(a) * sqrt(b)

OpenStudy (raden):

setting sqrt(96) = sqrt(16 * 6) now, apply formula above to get simplication form

OpenStudy (raden):

so, sqrt(96) = sqrt(16 * 6) = sqrt(16) * sqrt(6) = 4 sqrt(6)